OpenStudy (anonymous):

The weight of a piece of crystal in the form of a regular dodecahedron is 360 lbs., and each of its edges is 10 inches long. What is the weight of a similar piece of crystal whose edge edge measures 16 in.?

Directrix (directrix):

The two regular polyhedra of this problem are similar dodecahedrons. Theorem: The cube of the scale factor of two similar solids is equal to the ratio of the volumes of the two solids.

Directrix (directrix):

Therefore, (10/16) ^3 = 360/w where w is the weight of the dodecahedron with edge 16.

Directrix (directrix):

(10/16) ^3 = 360/w (5/8) ^ 3 = 360 / w 125/512 = 360/w Your task: solve for w @rossbhert_24